OpenStudy (anonymous):

CAN SOMEONE PPPLLLZZZ HELP ME!!!!!!!!!! what is the product of √t(5√3-4)

OpenStudy (anonymous):

is it \[\sqrt{t}(5\sqrt{3}-4)\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then there is not really much to do you could rewrite it as \[5\sqrt{3t}-4\sqrt{t}\] if you like by the distributive law

OpenStudy (anonymous):

ill give you the options hold on

OpenStudy (anonymous):

A.√2t B.√3t C.5√3t-4 D.5√3t-4√7

OpenStudy (anonymous):

hello?

OpenStudy (callisto):

He has told you the answer already